OpenStudy (anonymous):

In a parallelogram, a base b, and its corresponding height, h, are in the ratio of 5:3. The area is 135 mm squared. Find b and h.

OpenStudy (anonymous):

Since you have this ratio, you have: b/h = 5/3 which is b = 5h/3 Your area is b x h = 135 Substituting: (5h/3)(h) = 135 -> (5h^2)/3 = 135 Now, try solving for "h". Once you do that, you can use the original ratio to get "b".

OpenStudy (anonymous):

Thx for the recognition! @ajprincess

OpenStudy (anonymous):

How is this going for you, @samanthablank ?

OpenStudy (ajprincess):

u r welcome:) nd u really deserv it:)

OpenStudy (anonymous):

Those are worthy, kind words indeed. Especially coming from a 99'er.

OpenStudy (anonymous):

5h^2/ 3 = 135/1 5h^2=405 divide both sides by 5: h^2=81 square toot both sides: h=9 did I do that right? and now, I would just plug h into the ratios, right?

OpenStudy (anonymous):

You are doing very well here! Keep up the good work! Yes, just plug into the ratio now.

OpenStudy (anonymous):

ok, thanks! :)

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @samanthablank And you're very welcome!